OpenStudy (kainui):

I'm trying to solve this system of nonlinear equations and find the critical points. However, when I am solving for the critical points it seems that there are critical "lines" and I've never really heard of these before. x'=(2+x)(y-x) y'=(4-x)(y+x) So you can see when x'=0 and y'=0 I get: x=-2 x=4 y=x y=-x Should I just disregard these lines and just consider the points x=-2 and x=4 as my critical points? This seems wrong though.

OpenStudy (mathmale):

While not familiar with solving systems of nonlinear equations (which also happen to involve derivatives), I could point out a number of interesting things: first, y' = 0 when x=4 and/or when y=-x. Remember that y' generally signifies dy/dx and represents the slope of the tangent line to a curve at a given point. Are you assuming that y' =dy/dx? If so, what would x' =dx/dx signify?

OpenStudy (kainui):

No, x'=dx/dt and y'=dy/dt. They're independent variables in this case.

OpenStudy (anonymous):

Critical points would be t values in this case, right?

OpenStudy (kainui):

Not quite, I'm looking at the phase plane.

OpenStudy (anonymous):

Why do you want the critical point, and what exactly is it in this case? What variable?

OpenStudy (kainui):

Well in differential equations 2 there's a theorem that allows you to approximate a system of nonlinear differential equations as a linear system at those critical points. Essentially a linear system looks something like this: x'=y y'=-x And that will allow you to graph a vector field. Each vector is (x', y') located at (x,y).

OpenStudy (kainui):

In this example case I just made up, that will look something like this:

OpenStudy (kainui):

Well drawing doesn't work, but it's basically a bunch of concentric circles starting at the origin and increasing in magnitude as you move away.

OpenStudy (anonymous):

And what is the problem with the critical points you got?

OpenStudy (kainui):

Well y=x and y=-x aren't points, they're entire lines.

OpenStudy (anonymous):

For the system x'=y y'=-x What would you consider to be the critical points?

OpenStudy (kainui):

(0,0) since (x',y')=(0,0) when (x,y)=(0,0). But that was just an example explaining what a linear system is... My real question is in the first post.

OpenStudy (anonymous):

So you are saying that if for (x,y), then (x',y') = (0,0) then it is a critical point?

OpenStudy (kainui):

Exactly.

OpenStudy (anonymous):

Okay, first start with x = -2 This makes x' = 0, Now we need to make y' = 2. We have y' = 0 = (4-(-2))(y-2) solve for y

OpenStudy (mathmale):

@Kainui : No, x'=dx/dt and y'=dy/dt. They're independent variables in this case. Thanks for the clarification! Mind if I make small corrections here? x and y are the dependent variables and t, the parameter, is the independent variable. :)

OpenStudy (anonymous):

Now we need to make y' = 0 ^^

OpenStudy (anonymous):

So one possible critical point would be (x, y) = (-2, 2)

OpenStudy (anonymous):

Do the same with x=4 Then consider x-y = 0 = x+y

OpenStudy (kainui):

@mathmale Well since x and y are the basis of a vector space, they are absolutely linearly independent. In fact, they don't exactly depend on t in the way you're thinking since the phase plane represents the infinite number of solutions that satisfy the differential equation. @wio That's exactly what I've done in the very first post, and now you're where I am I believe. Since y=x, I can easily have the following critical points: (5,5), (7,7), (1000,1000),... So... Yeah.

OpenStudy (anonymous):

But reconsider (5,5) x'=(2+5)(5-5) = 0 y'=(4-5)(5+5) = (-1)(10) = -10

OpenStudy (anonymous):

I'm not saying systems with infinite critical points don't exist, but I'm not sure this is such a system.

OpenStudy (kainui):

Ahhh, duh, it seems pretty obvious now what you're saying, let me work it out real quick thanks!

OpenStudy (anonymous):

Unless (x',y') = (0,c) or (x',y') = (c,0) are critical too

OpenStudy (kainui):

Ok just to check, I got three critical points, (-2,2),(4,4) and(0,0). Does that seem right to you?

OpenStudy (anonymous):

Yes

OpenStudy (kainui):

Cool. Thanks.

OpenStudy (anonymous):

Profiling this. :3